# Magnetocalorics of singlet ground state induced moment magnets

### v1

@@ Appendix A Adiabatic caloric coefficients for noninteracting two singlet model
As a reference to the FM and AFM models it is useful to derive the adiabatic caloric coefficients for the case of isolated two-level systems. Using their definitions in Sec.  IV.2 and setting $I_{e}$= 0 or $\xi_{s}$= 0 in the thermodynamic potentials we can derive the expressions (for clarity of dimensions we explicitly include kB)
$\Gamma^{0}_{mc}(T,H)=\bigl{(}\frac{k_{B}T}{\Delta_{h}}\bigr{)}\bigl{(}\frac{m_{s}h}{\Delta_{h}}\bigr{)}\;\;\bigl{[}\frac{g_{eff}\mu_{B}}{k_{B}}\bigr{]}$ (52)
where $\Delta_{h}$=$[\Delta^{2}+(m_{s}h)^{2}]^{\frac{1}{2}}$ is the two singlet splitting in an external field but without interaction effect. This expression is proportional to temperature as well as field strength. This is quite different from the result for a degenerate magnetic Kramers doublet which is, following a similar procedure, given by
$\Gamma^{0}_{mc}(T,H)=\bigl{(}\frac{k_{B}T}{m_{s}h}\bigr{)}\;\;\bigl{[}\frac{g_{eff}\mu_{B}}{k_{B}}\bigr{]}=\frac{T}{H}$ (53)
@@
$\Gamma^{0}_{bc}(T,p)=\Bigl{(}\frac{k_{B}T}{\Delta(p)}\Bigr{)}\Bigl{(}\frac{\Delta(p)}{\Delta_{h}(p)}\Bigr{)}^{3}k_{B}^{-1}$ (54)
Here we implied that the pressure in $(\partial T/\partial p)_{S}$ is given in units of $[c_{B}/\Omega_{\Delta}]$. In these units the pressure dependence of the bare singlet splitting is given by $\Delta(p)=\Delta(1+p)$ in the Grüneisen model. Without applied field this reduces to
$\Gamma^{0}_{bc}(T,p)=\Bigl{(}\frac{k_{B}T}{\Delta(p)}\Bigr{)}k_{B}^{-1}$ (55)

### v2

@@ Appendix A Adiabatic caloric coefficients for noninteracting two singlet model
As a reference to the FM and AFM models it is useful to derive the adiabatic caloric coefficients for the case of isolated two-level systems. Using their definitions in Sec.  IV.2 and setting $I_{e}$= 0 or $\xi_{s}$= 0 in the thermodynamic potentials we can derive the following expressions (for clarity of dimensions we explicitly include kB):
$\Gamma^{0}_{mc}(T,H)=\bigl{(}\frac{k_{B}T}{\Delta_{h}}\bigr{)}\bigl{(}\frac{m_{s}h}{\Delta_{h}}\bigr{)}\;\;\bigl{[}\frac{g_{eff}\mu_{B}}{k_{B}}\bigr{]},$ (52)
where $\Delta_{h}$$\Delta^{2}+(m_{s}h)^{2}]^{\frac{1}{2}}$ is the two singlet splitting in an external field but without interaction effect. This expression is proportional to temperature as well as field strength. This is quite different from the result for a degenerate magnetic Kramers doublet that is, following a similar procedure, given by
$\Gamma^{0}_{mc}(T,H)=\bigl{(}\frac{k_{B}T}{m_{s}h}\bigr{)}\;\;\bigl{[}\frac{g_{eff}\mu_{B}}{k_{B}}\bigr{]}=\frac{T}{H},$ (53)
@@
$\Gamma^{0}_{bc}(T,p)=\Bigl{(}\frac{k_{B}T}{\Delta(p)}\Bigr{)}\Bigl{(}\frac{\Delta(p)}{\Delta_{h}(p)}\Bigr{)}^{3}k_{B}^{-1},$ (54)
with $\Delta_{h}(p)$$\Delta(p)^{2}+(m_{s}h)^{2}]^{\frac{1}{2}}$. Here we implied that the pressure in $(\partial T/\partial p)_{S}$ is given in units of $c_{B}/\Omega_{\Delta}]$. In these units the pressure dependence of the bare singlet splitting is given by $\Delta(p)=\Delta(1+p)$ in the Grüneisen model. Without applied field this reduces to
$\Gamma^{0}_{bc}(T,p)=\Bigl{(}\frac{k_{B}T}{\Delta(p)}\Bigr{)}k_{B}^{-1},$ (55)